OpenStudy (jpes2193):

x^2/3-3x^1/3-4=0 Solve by making a u substitution.

OpenStudy (freckles):

let u=x^(1/3) so u^2=x^(2/3) you have a quadratic equation in terms of x^(1/3) not in terms of u since you replaced all the x^(1/3) 's with u

OpenStudy (jpes2193):

okay

OpenStudy (jpes2193):

u^2-3u-4=0

OpenStudy (freckles):

ok solve for u factor or use quadratic formula